# Quantum mechanics provides the physical basis of teleological evolutions

### v1

@@ Abstract
We show that the quantum computational speedup of quantum algorithms is due to their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Naturally, this is a destabilizing result: the teleological notion has long been rejected by science and even, for its finalistic character, viewed as a belief. We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living for which the teleological notion was conceived. This work should show that the discovery of the quantum computational speedup confers full scientific dignity to the holistic and finalistic view of physical reality traditionally opposed by science.
@@ 1 Introduction
According to natural philosophy, teleological evolutions are evolutions attracted by a final cause; obviously, it would be causation from the future, i.e. retrocausation. Before the advent of modern science, the teleological notion was applied to the evolutions of life, where it provided a potentially enlightening relation between evolution and functional structure of the living: teeth were created for the purpose of chewing. As is well known, with the advent of modern science the teleological notion was dismissed for the alleged absence of a physical basis; for its finalistic character, it was also ascribed to beliefs.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup [^1] By quantum (computational) speedup, we will always mean its certain form, the fact that a quantum algorithm certainly solves its problem with fewer logical operations (computation steps) than possible in the classical world. should change this position of science. This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since this will turn out to be surprisingly evident, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to a deeply rooted rejection, on the part of physicists, of the related notions of retrocausality and finalism. The notion of teleological evolution, of course, involves both.
We, with coauthors, already went against this possible refusal in the evolutionary approach to quantum speedup [1, 2, 3, 4, 5, 6]. In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obey the following rule:
@@
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem); let us underline the novelty of this concept. Under the advanced knowledge rule, in addition to responding to the philosophical notion of teleological evolution, quantum algorithms obviously provide a scientific example of it.
@@
Until now, the advanced knowledge rule was, so to speak, an empirical rule satisfied by all the major quantum algorithms. In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way. Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
@@
We now go to the evolutions of living to which the teleological notion was first applied. Naturally, one could see an enormous difference between quantum algorithms and the evolutions of living. The former have maximum durations of a few dozen microseconds and a rather limited number of quantum bits, on the order of one thousand; beyond that, today, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, interestingly, the present context suggests a way out. The difference is enormous in the usual classical way of viewing macroscopic reality but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such (as we shall see), it can occur with quantum speedup and thus be teleological in character.
To see that the evolutions of living have the form of quantum algorithms, in the first place, we should extend the usual quantum description of quantum algorithms, which is limited to the unitary computation of the measurement of the solution, to the upstream process of setting the problem. We also need some terminology. Consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers; we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. Of course, the problem-setting (the number of the drawer with the ball) is here also the solution of the problem. The extended quantum description of the quantum algorithms is thus as follows: an initial measurement randomly selects the problem-setting out of a quantum superposition (or, indifferently, mixture) of all the possible problem settings, the unitary computation of the corresponding solution builds up one-to-one correlation between the problem-setting selected by the initial measurement and the solution, the final measurement reads the solution.
Note that, according to the advanced knowledge rule, in the optimal quantum case, the number of computation steps needed by the quantum algorithm is that needed in the optimal logical case benefiting from the advanced knowledge of half of the information about the solution. Also note that, however, this is just the room for the quantum speedup. The quantum algorithm can fully benefit of it, or partly, or not at all.
@@
Now we go to the evolution of the universe under the fundamental physical laws and the fundamental constants. The version of the Anthropic Principle we are dealing with is about the fact that the least change of the value (for values) that the fundamental physical constants have in the universe we are in would have produced a relatively trivial universe unable to develop life. This can be formulated as follows.
We should think of a plurality of universes, all with the same fundamental laws and each with its own value of the fundamental constants. We consider any one of these universes, in the mutually exclusive or of all the possible universes.
The Fine-tuned Universe version of the Anthropic Principle says that its evolution under the same fundamental physical laws and its own value of the physical constants, builds up a (say) one-to-one correlation between the value of the fundamental constants since the beginning of time and the “value of life” (in particular, “existence of sentient life” and “absence of life”).
Under the quantum cosmological assumption, the evolution of the universe toward the value of life becomes unitary quantum and the one-to-one correlation becomes quantum. Under the quantum cosmological reading, the plurality of universes becomes their quantum superposition and the selection of any one universe in the mutually exclusive or of all universes can be performed by an initial measurement (see what follows) of the value of the fundamental constants in the quantum superposition of all the possible universes.
This is formally identical to a quantum algorithm (see the above extended quantum description of it). Therefore, as we shall see, it can occur with quantum speedup, in the sense that there is certainly the room for its occurring with quantum speedup. Correspondingly, it can be teleological in character.
By the way, the notion of measurement requires that of an observer, and, no matter the value of the fundamental constants, at the beginning of time there could have been none. However, we can do without this initial measurement. There would then be a unitary evolution of the quantum superposition of all possible universes. In the universe we are in, and at the time of today, there would be the sentient observer who measures the value of life also selecting the corresponding value of the fundamental constants. According to the version of the Anthropic Principle we are dealing with, at the time of today these two observables would already be maximally entangled. Since the value of the fundamental constants never changes along the unitary evolution of the universe, all is as if it were selected at the beginning of time.
@@
Summarizing, in the universe we are in, its evolution toward life can have occurred with quantum speedup (in the sense that there is certainly the room for this) and thus have been teleological in character. We shall also see that the same would consequently hold for the evolution of life on Earth (in principle, the trace over all the rest of that of the universe toward life) and the evolutions of species.
By the way, all the above, the quantum cosmological assumption and the consequent reading of the Fine-tuned Universe version of the Anthropic Principle, can, of course, be considered arbitrary. We would argue that it is anyway a plausible possibility, thus very different from the current notion that a teleological evolution of life is an absurdity.
Of course, if taken seriously, the above would also be highly destabilizing. Besides providing a unified quantitative explanation of the quantum speedups, the present result would change the very foundations of the natural sciences, restoring what could be an enlightening relation between evolution and functional structure of the living. Again, our purpose here is showing that quantum algorithms are certainly teleological evolution, and their distance from the evolutions of life could even be zeroed under a plausible way of viewing macroscopic reality, i.e. the quantum cosmological assumption.
Also note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they could acquire a physical basis.
@@ 2 Time-symmetric explanation of the quantum speedup
The present way of thinking about quantum algorithms and their quantum speedups has radical differences from the mainstream one. In this section, we provide it at a conceptual level, but always making a point-by-point reference to the formalizations, all postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that the discovery of the quantum speedup has been one of the fundamental discoveries of the past century, which, together with quantum communication, has created an entirely new branch of physics. However, having said this, for the sake of clarity, we will be completely frank in criticizing what we consider to be the shortcomings of the current way of viewing quantum computation.
2.1 On the explanation of the quantum speedup
Let us begin with the advanced knowledge rule, which we repeat here for convenience.
@@ 2.1 On the explanation of the quantum speedup
Of course, this rule offers a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, the absence of such an explanation has never been acknowledged by the scientific community.
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, cannot be considered its explanation. By the way, what distinguishes science from opinion is, in particular, its quantitative character. The fact that the most popular explanation of the quantum speedup, a quantitative feature, is qualitative in character, would seem to be rather unscientific. The only justification we can see is that this is a way of removing the fact that, forty years after its discovery—an explanation of the quantum speedup is still missing. This is naturally our opinion. However, as a matter of fact, outside our (destabilizing) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
@@
In our view, for what concerns the current mainstream explanation of the quantum speedup, this is the case of crying out: the emperor is naked! Our emphasis on this is because it could balance the high unconventionality of our proposal.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against a deeply rooted de facto ban on retrocausality. In this work, we should have demonstrated it, indeed in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the destabilizing character of the rule.
@@ 2.2 First demonstration of the advanced knoledge rule
In any case, in order to demonstrate the advanced knowledge rule, we should first amend the current quantum description of quantum algorithms: as it is, it would completely obscure both demonstrations. The quantum description in question is limited to the unitary computation of the solution, where the quantum speedup resides, and the measurement of the latter. Of course, a canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. The current quantum description is obviously focused on the central problem of how to achieve a quantum speedup with respect to to what is logically possible. By the way, the explanation of the quantum speedup we will be dealing with completely ignores the way in which it is reached, it is limited to the fact that it is reached.
Anyway, the above shortcoming is simply remedied: we should add to the current quantum description of the process of solving the problem that of the upstream process of setting the problem. An initial measurement in the quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random [^2] The problem-setter could unitarily change it at will; this, here, would change nothing., the unitary computation and the final measurement of the corresponding solution follow—for the formalization of this, see Section (1) of the Appendix. We would recommend to look at it; it is trivial but introduces all the definitions that will be used in the subsequent formalizations.
The quantum description of the quantum algorithm defined until now is with respect to the customary observer external to the quantum process. To present ends, we should relativize it with respect to the problem-solver, who computes the solution of the problem. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her. From her, the problem-setting selected by the initial measurement should be concealed. As we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action. This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”. It is ordinary because it is still the customary “causal” description of mainstream quantum mechanics—where causality always goes in the same time-direction.
The first demonstration of the advanced knowledge rule simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
@@
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [7]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [20] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference.
We show how a classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule. For the moment, we consider the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes to the full fledged Grover algorithm and most general forms of quantum correlation (in quantum algorithms, between the problem-setting and the solution of the problem).
@@
Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings). There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting of this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm.
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert in which the search for the solution occurs reduces from $N$ to $\sqrt{N}$.
@@
We shall see further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.

### v2

@@ Abstract
We show that the quantum computational speedup is due to the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
@@ 1 Introduction
Let us briefly examine the very long history of the teleological notion—attraction by a final cause. Plato and Aristotle used it to explain the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution. With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy. It was eventually dismissed by science for the alleged absence of a physical basis. All this was before the advent of quantum mechanics. Today, the situation of the teleological notion is somewhat complex. To the best of our knowledge, outside the present work, there are no physical researches on it. Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science. This is because, in the first place, quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since showing this will be surprisingly simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
We, with coauthors, already went against this position in the evolutionary, retrocausal, approach to the quantum computational speedup [1, 2, 3, 4, 5, 6]. In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obeyed the following rule:
@@
By the way, here and in the following by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible. Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
@@
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem).
@@
Until now, this advanced knowledge rule was, so to speak, an empirical rule shown to be satisfied by all major quantum algorithms. In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way. Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
As a matter of fact, we have provided two redundant demonstrations of this rule, what should not be despised given its destabilizing character. Let us outline them.
One is an application of classical logic to what the quantum algorithms does: providing a quantum speedup with respect to what is classically possible.
By the way, we will have to use a quantum description of the quantum algorithms that is not their customary quantum description. The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement. We complete the customary quantum description by adding to the description of the process of solving the problem that of the upstream process of setting the problem. The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow. Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic. As we shall see in the next section, the classical logic description of the quantum speedup is indeed the advanced knowledge rule.
Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete. Indeed, it is simply completed by time-symmetrizing it. We shall see that the time-symmetrization of the quantum description yields the quantum superposition of the causal loops involved in the advanced knowledge rule. In other words, the time-symmetrization of the quantum description of the quantum algorithm generates the physical correspondent of the advanced knowledge rule.
The second demonstration of the advanced knowledge rule simply comes from the fact that this time-symmetrization of the quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far. As we shall see, it is to satisfy an atemporal symmetry implicit in the quantum description. Of course, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—it is a demonstration of the rule.
Of course, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
A further step, important for what concerns the possible teleological character of the evolutions of the living, is that all the above generalizes to quantum correlation. For simplicity, we limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty. We can then say that the quantum algorithm builds up one-to-one correlation between two initially unrelated observables (the problem-setting and the solution of the problem) without ever changing the eigenvalue of the observable selected by the initial measurement (the problem-setting). We shall show that any building up of quantum correlation of this kind (not only that in quantum algorithms) can occur with quantum speedup and therefore be teleological in character. Note that the “can be” means that there is certainly room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation. By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
With this, we can go to the evolutions of the living for which the teleological notion was originally conceived.
First, let us note that the dismissal of the teleological notion was not without (scientific) regret; indeed, it meant doing without a possibly enlightening relation between the notion of evolution of life and that of the functional structure of the living. We cite two authoritative objections to this dismissal. One is a famous sentence by the evolutionary biologist Haldane [8]:
Let us now go to what the teleological character of quantum algorithms can say to the evolutions of the living.
@@
Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living. The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, interestingly, the very context we are dealing with suggests a way out. The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11]. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such, it can occur with quantum speedup and thus be teleological in character.
Let us describe how things should go.
@@
The Fine-tuned universe version of the Anthropic Principle is about the very relevant fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
Under the quantum cosmological assumption, this version of the Principle can be formalized as follows. We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
Now, always under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental physical laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental physical constants, and each with its own value of the latter. Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”. The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life. Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life; in the other universes, where this value is different, there is absence of life (theory says that, the absence of life in the other universes today is already definitive).
For what we have said about the generalization of the quantum speedup from quantum algorithms to the building up of quantum correlation, the evolution of the universe toward the value of life can occur with quantum speedup and, therefore, be teleological in character. The “can occur” is in the sense that there is room for its occurring with quantum speedup.
@@
Of course, in the universe we are in, the possible quantum speedup and the consequent teleological character is for the evolution of the universe toward life. We will see that this possibility would be inherited by the evolution of life on Earth, in principle, the trace over all the rest of the evolution of the universe toward life. Then, for consistency, it would be inherited by the evolutions of species.
Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered as plausible as the quantum cosmological assumption. However, note that all this is quite different from the present widespread belief that the notion of teleological evolution of the living is an absurdity.
Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they would acquire a physical basis. Of course, this would not touch on the immense practical value of the mechanistic and reductionist view. That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe obviously holds only in principle and does not give what the classical description gives. However, this might pinpoint a direction of the research. At the theoretical level, it could be interesting to study how this “tracing over all the rest” conciliates a holistic and teleological view in the longer term with a mechanistic and reductionist view in the shorter term. This might even be a further perspective under which to explore the unification between classical and quantum mechanics.
@@ 2 Time-symmetric explanation of the quantum speedup
We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality. Since this is a way of thinking quantum computation radically different from the usual one, we provide it first at a conceptual level. All formalizations are postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century. Together with that of quantum communication, it has created an entirely new branch of physics. This said, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
2.1 On the explanation of the quantum computational speedup
We repeat here for convenience the advanced knowledge rule.
@@ 2.1 On the explanation of the quantum computational speedup
Of course, this rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, the absence of a unified quantitative explanation has never been acknowledged by the scientific community.
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, cannot be considered its explanation. By the way, what distinguishes science from opinion is, in particular, its quantitative character. The fact that the most popular explanation of the quantum speedup, a quantitative feature, is qualitative in character, would seem to be rather unscientific. The only justification we can see is that this is a way of removing the fact that, forty years after its discovery—an explanation of the quantum speedup is still missing. This is naturally our opinion. However, as a matter of fact, outside our (unconventional) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
Of course, as already noted, if the only way of providing the unified quantitative explanation in question is by resorting to retrocausality and finalism, the de facto rejection of these notions by the scientific community might have played a role.
@@
For what concerns the current mainstream explanation of the quantum speedup, let us say, it is the case of crying out that the emperor is naked. This emphasis is to underline the absence of any recognition of the fact that the current view of quantum computation has a shortcoming.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against the deeply rooted de facto ban on retrocausality. In this work, we should have demonstrated it, moreover, in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the unconventional character of the rule.
@@ 2.2 First demonstration of the advanced knoledge rule
We have already given in the Inroduction the quantum description of quantum algorithms extended to the process of solving the problem. An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing.. The unitary computation and the final measurement of the corresponding solution follow. For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
Now, it should also be noted that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. To present ends, we should also relativize it with respect to the problem-solver, who computes the solution of the problem. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her. Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action. Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easily obtained. It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of the unitary evolution.
This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”. It is ordinary because it is still the customary “causal” quantum description of mainstream quantum mechanics—where causality always goes in the same time-direction.
We can go now to the first demonstration of the advanced knowledge rule. It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
@@
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [12]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [13] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference. By the way, note that the notion of searching for something and finding it is most universal and this universality is completely represented by Grover algorithm.
We show that the classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule. We begin with the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
@@
Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way, one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings); the fewest is the number needed in the optimal logical case benefiting of half of the information about the solution. There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting from this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm.
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
@@
We shall see that there are further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.